If (2/5)x + (1/3)y =1, then what is. 12x +10y?
Answer by Alan3354(69443)

If (2/5)x + (1/3)y =1, then what is. 12x +10y?
---------
(2/5)x + (1/3)y =1
Times 15
6x + 5y = 15
Times 2
12x + 10y = 30